Superior Micro Products uses the weighted-average method in its process costing system. During January, the Delta Assembly Department completed its processing of 25,400 units and transferred them to the next department. The cost of beginning work in process inventory and the costs added during January amounted to $690,904 in total. The ending work in process inventory in January consisted of 2,800 units, which were 70% complete with respect to materials and 50% complete with respect to labor and overhead. The costs per equivalent unit for the month were as follows: Materials Labor Overhead $13.40 4.50 7.60 Cost per equivalent unit Required: 1. Compute the equivalent units of materials, labor, and overhead in the ending work in process inventory for the month. 2. Compute the cost of ending work in process inventory for materials, labor, overhead, and in total for January. 3. Compute the cost of the units transferred to the next department for materials, labor, overhead, and in total for January. 4. Prepare a cost reconciliation for January. (Note: You will not be able to break the cost to be accounted for into the cost of beginning work in process inventory and costs added during the month.)Complete this question by entering your answers in the tabs below. Required1Required 2Required 3 Required 4 Compute the equivalent units of materials, labor, and overhead in the ending work in process inventory for the month. Materials Labor Overhead Equivalent units Required1 Required 2>Complete this question by entering your answers in the tabs below. Required 1 Required 2Required 3 Required 4 Compute the cost of ending work in process inventory for materials, labor, overhead, and in total for January. Materials Labor Overhead Total Cost of ending work in process inventory K Required 1 Required 3>Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3Required 4 Compute the cost of the units transferred to the next department for materials, labor, overhead, and in total for January Materials Labor Overhead Total Cost of units completed and transferred out Required 2 Required 4Complete this question by entering your answers in the tabs below. Required 1Required 2 Required 3 Required 4 Prepare a cost reconciliation for January. Cost Reconciliation Total cost to be accounted for Costs accounted for as follows: Total cost accounted for Required 3 Required 4
Answer #1

Calculate equivalent unit

Material Labor Overhead
Unit Completed and transferred out 25400 25400 25400
Ending work in process unit 2800*70% = 1960 2800*50% = 1400 1400
Equivalent unit of production 27360 26800 26800

Cost of ending work in process

Material Labor Overhead Total
Cost of ending work in process 1960*13.40 = 26264 1400*4.50 = 6300 1400*7.60 = 10640 43204

Cost of unit completed and transferred out

Material labor Overhead Total
Cost of unit completed and transferred out 25400*13.40 = 340360 25400*4.50 = 114300 25400*7.60 = 193040 647700

Cost reconciliation

Total Cost to be accounted for 690904
Cost accounted for as follows
Cost of unit completed and transferred out 647700
Cost of ending work in process 43204
Total Cost accounted for 690904